OpenStudy (bossimbacon):

Francine and Cheryl received equal scores on a test made up of multiple choice questions and an essay. Francine got 34 multiple choice questions correct and received 14 points for her essay. Cheryl got 30 multiple choice questions correct and received 22 points for her essay. How many points was each multiple choice question worth? ___ points

OpenStudy (anonymous):

34x+14 = 30x+22 solve for x

OpenStudy (bossimbacon):

2

OpenStudy (anonymous):

x=2 yea

OpenStudy (bossimbacon):

so whats the answer

OpenStudy (anonymous):

2 your answer

OpenStudy (bossimbacon):

In May, Liam and Charlie had the same amount of money in their savings accounts. In June, Liam deposited $140 into his account. Charlie said he increased the money in his account by 5%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in May? $

OpenStudy (bossimbacon):

In May, Liam and Charlie had the same amount of money in their savings accounts. In June, Liam deposited $140 into his account. Charlie said he increased the money in his account by 5%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in May? ___$

OpenStudy (anonymous):

amt of money Liam = a Charlie = b a=b a + 140 b + 0.05b a+140 = a + 0.05a 140 = 0.05a 2800 = a

OpenStudy (bossimbacon):

2800 is the answer

OpenStudy (anonymous):

yes